# An innovative circuit for testing hot carrier and trap generation in GaN Devices

### v1

@@ III-B Contact Degradation
Contact degradation and gate sinking are significant degradation mechanisms at elevated temperatures in GaAs and InP based HEMTs. This has not yet proven to be a significant issue with AlGaN/GaN HEMTs at temperatures below 400 °C for Pt/Au Schottky contacts and Ti/Al/Pt/Au annealed Ohmic contacts [Meneghini2016PowerGaNDevices, Uemoto2007GateInjectionTransistor, Rossetto2016StudyOfTheStability, Meneghini2015ExtensiveInvestigation, Degraeve1998NewInsights, Stathis1999PercolationModels, Meneghini2016NormallyOffGaN]. An increase in Schottky barrier height was observed for Ni/Au Schottky contacts after DC stress at elevated junction temperatures (200 °C) [Meneghini2016PowerGaNDevices, Uemoto2007GateInjectionTransistor, Rossetto2015DemonstrationOfField, Chen2012HBMEsdRobustness, Shankar2016UniqueEsdBehavior, Kuzmik2004ElectricalOverstress]. This was due to consumption of an interfacial layer between the Schottky contact and the AlGaN layer. Though the resulting positive shift in the Schottky barrier height, and thus the pinch-off voltage, is ideal, the subsequent change in IDSS is not favorable. Unstressed devices were subjected to an anneal after the Schottky contact was deposited in order to decrease the interfacial layer between the gate and semiconductor. Devices that underwent the gate anneal showed 50% less degradation during a 24 h stress test as opposed to devices that did not receive a gate anneal [Meneghini2016PowerGaNDevices, Uemoto2007GateInjectionTransistor]. Thermal storage tests up to 2000 h on Ti/Al/Ni/Au ohmic contacts at and above 290 °C showed an increase in contact resistance as well as surface roughness due to growth of Au-rich grains that ultimately led to cracks in passivation [Rossetto2015DemonstrationOfField, Chen2012HBMEsdRobustness, Shankar2016UniqueEsdBehavior, Khan1994CurrentVoltageCharacteristic, Vetury2001ImpactOfSurfaceStates, Hu2001Si3N4AlGaN, Leirer2001RFCharacterization, Kuzmik2004ElectricalOverstress]. The two primary degradation mechanisms were determined to be Au inter-diffusion within the metal layers and Ga out-diffusion from the semiconductor into the metallic compounds. Similar degradation was observed after DC stress tests that resulted in junction temperatures equivalent to the thermal storage tests. Due to the high-power capability of AlGaN/GaN HEMTs, proper temperature management is needed in order to optimize device performance under high current and high voltage operation [Kim2003EffectsOfSiNPassivation, Ibbetson2000PolarizationEffects, Gao2014ImpactOfWaterAssisted, Ambacher2002PyroelectricProperties, Joh2008CriticalVoltage, valizadeh2005effects]. Self-heating of devices can ultimately result in poor device performance through contact degradation. Reliability of contacts is highly dependent upon both metal schemes as well as processing during fabrication, EPC’s eGaN transistors that were used in this research do not use a Schottky gate and therefore are not vulnerable to this whole contact mechanism.

### v2

@@ III-B Contact Degradation
Contact degradation and gate sinking are significant degradation mechanisms at elevated temperatures in GaAs and InP based HEMTs. This has not yet proven to be a significant issue with AlGaN/GaN HEMTs at temperatures below 400 °C for Pt/Au Schottky contacts and Ti/Al/Pt/Au annealed Ohmic contacts [Meneghini2016PowerGaNDevices, Uemoto2007GateInjectionTransistor, Rossetto2016StudyOfTheStability, Meneghini2015ExtensiveInvestigation, Degraeve1998NewInsights, Stathis1999PercolationModels, Meneghini2016NormallyOffGaN, inberg2001material]. An increase in Schottky barrier height was observed for Ni/Au Schottky contacts after DC stress at elevated junction temperatures (200 °C) [Meneghini2016PowerGaNDevices, Uemoto2007GateInjectionTransistor, Rossetto2015DemonstrationOfField, Chen2012HBMEsdRobustness, Shankar2016UniqueEsdBehavior, Kuzmik2004ElectricalOverstress, axelevitch2014application]. This was due to consumption of an interfacial layer between the Schottky contact and the AlGaN layer. Though the resulting positive shift in the Schottky barrier height, and thus the pinch-off voltage, is ideal, the subsequent change in IDSS is not favorable. Unstressed devices were subjected to an anneal after the Schottky contact was deposited in order to decrease the interfacial layer between the gate and semiconductor. Devices that underwent the gate anneal showed 50% less degradation during a 24 h stress test as opposed to devices that did not receive a gate anneal [Meneghini2016PowerGaNDevices, Uemoto2007GateInjectionTransistor]. Thermal storage tests up to 2000 h on Ti/Al/Ni/Au ohmic contacts at and above 290 °C showed an increase in contact resistance as well as surface roughness due to growth of Au-rich grains that ultimately led to cracks in passivation [Rossetto2015DemonstrationOfField, Chen2012HBMEsdRobustness, Shankar2016UniqueEsdBehavior, Khan1994CurrentVoltageCharacteristic, Vetury2001ImpactOfSurfaceStates, Hu2001Si3N4AlGaN, Leirer2001RFCharacterization, Kuzmik2004ElectricalOverstress, golan2000novel]. The two primary degradation mechanisms were determined to be Au inter-diffusion within the metal layers and Ga out-diffusion from the semiconductor into the metallic compounds. Similar degradation was observed after DC stress tests that resulted in junction temperatures equivalent to the thermal storage tests. Due to the high-power capability of AlGaN/GaN HEMTs, proper temperature management is needed in order to optimize device performance under high current and high voltage operation [Kim2003EffectsOfSiNPassivation, Ibbetson2000PolarizationEffects, Gao2014ImpactOfWaterAssisted, Ambacher2002PyroelectricProperties, Joh2008CriticalVoltage, valizadeh2005effects]. Self-heating of devices can ultimately result in poor device performance through contact degradation. Reliability of contacts is highly dependent upon both metal schemes as well as processing during fabrication, EPC’s eGaN transistors that were used in this research do not use a Schottky gate and therefore are not vulnerable to this whole contact mechanism.